### v1

# Devising PoPStat: A metric bridging population pyramids with global disease mortality

Abstract
Understanding the relationship between population dynamics and disease-specific mortality is central to evidence-based health policy. This study introduces two novel metrics, PoPDivergence and PoPStat, one to quantify the difference between population pyramids and the other to assess the strength and nature of their association with the mortality of a given disease. PoPDivergence, based on Kullback-Leibler divergence, measures deviations between a country’s population pyramid and a reference pyramid. PoPStat is the correlation between these deviations and the log form of disease-specific mortality rates. The reference population is selected by a brute-force optimization that maximizes this correlation. Utilizing mortality data from the Global Burden of Disease 2021 and population statistics from the United Nations, we applied these metrics to 371 diseases across 204 countries. Results reveal that PoPStat outperforms traditional indicators such as median age, GDP per capita, and Human Development Index in explaining the mortality of most diseases. Noncommunicable diseases (NCDs) like neurological disorders and cancers, communicable diseases (CDs) like neglected tropical diseases, and maternal and neonatal diseases were tightly bound to the underlying demographic attributes whereas NCDs like diabetes, CDs like respiratory infections and injuries including self-harm and interpersonal violence were weakly associated with population pyramid shapes. Notably, except for diabetes, the NCD mortality burden was shared by constrictive population pyramids, while mortality of communicable diseases, maternal and neonatal causes and injuries were largely borne by expansive pyramids. Therefore, PoPStat provides insights into demographic determinants of health and empirical support for models on epidemiological transition. By condensing the multi-dimensional population pyramid to a scalar variable, PoPDivergence allows us to examine the relationship of any variable of interest with the underlying population structure. These metrics have the potential to advance future global health and population research.
1Department of Electrical and Electronic Engineering, Faculty of Engineering, University of Peradeniya.
2Department of Community Medicine,3Department of Pharmacology,
Faculty of Medicine, University of Peradeniya.
1 Introduction
Understanding the social and economic determinants of mortality causes is crucial for developing effective healthcare policies, interventions, and prevention programs. To this end, many studies have explored factors that help explain the variation in mortality causes. For instance, the Human Development Index (HDI), a composite metric combining life expectancy, education, and income, is used to examine cancer mortality patterns across regions [1]. These factors are often challenging to measure directly due to underlying, unobserved circumstances, commonly referred to as latent variables. Latent variables have been known to impact multiple causes of death. Among them, demographic variables are particularly significant, as they can provide valuable insights into observed mortality trends.
The effects of demographic and cultural variations, as well as how different groups responded to various diseases, have been examined using factors such as median age, income level and population density, all of which are valuable for policymaking[2, 3, 4]. Furthermore, demographic parameters have been utilized to track disease progression within communities, and correlating these parameters with disease spread has become a common practice for both communicable[5, 6, 7, 8] and non-communicable[9, 10, 11] diseases. A recent study by Carnegie et al indicates that high population density appears to be associated with high mortality rates of various non-communicable diseases (NCD) in western developed countries[12]. Additionally, research by Li et al. demonstrates that epidemic sizes are strongly linked to population density[13], while a study by Pardhan et al. shows a negative relationship between GDP per capita and the increase in new COVID-19 cases across European countries. These scalar variables, particularly the median age, are derived from a country’s population distribution. However, due to the median being an average-like statistic, it does not provide insights into the spread or shape of the distribution, which is essential for understanding the population structures. Consequently, the characteristics associated with different population pyramid shapes, expansive and constrictive forms, are not captured by the median age, which fails to reflect demographic variation. Regardless, the vast usage of median age proves the applicability of the information from the population distribution even in its most basic form is warranted[14, 15, 16].
The population pyramid breaks down the population according to gender: male, female, and age group. As shown in figure 1, the most prominent population pyramid structures are expansive, stationary, and constrictive. Expansive structures, also referred to as progressive, are considered youthful populations with a broad base indicating high birth rates. Stationary structures represent populations that are stable in size, with birth and death rates approximately balancing each other, leading to a roughly uniform distribution across age groups. Constructive structures, also referred to as regressive, are considered aged populations with a narrower base indicating low birth rates. Nevertheless, for brevity, these quantitative data of each country will be referred to as the population pyramid[17].
(a) Expansive
(b) Stationary
(c) Constrictive
Figure 1: Prominent population pyramid structures
The evolution of the pyramid shape is a result of historical events[18] such as major policy changes, natural phenomena, and political events as these factors affect the birth and death rate of countries. For example, the Middle East population doubled from 25 million to 54 million between 2005 and 2015 as the number of migrant individuals and families increased due to economic opportunities and China’s pyramid shrank inwards for the age cohorts around 50, showing the slowing of population growth during the Great Chinese famine. After the famine, the population started increasing, but about 30 years ago, the pyramid started thinner again with the “one-child policy”[19, 20] enacted in 1979. Hence, the population pyramid reflects national policies[18] and how said policies affect the country’s living conditions.
Apart from the derivative variables of the population pyramid, such as median age, a correlation exists between the economic variables and the pyramid shape[21]. For example, a macroeconomist can analyze the pyramid shape of a country to explore the relationship between population growth and GDP per capita growth. Similarly, a health economist might examine how local population composition impacts the demand for long-term care facilities. Besides, countries with different population compositions (but similar in other respects) will allocate their resources differently. For example, an inverted pyramid implies the government could negotiate more health service payments. Consequently, their GDPs will also differ, as countries with stronger government control over healthcare costs may allocate resources differently, leading to varying economic outcomes. Understanding these connections is crucial for policymakers seeking to address the specific needs of a population based on its demographic structure.
These economic and societal correlations highlight the population pyramid’s potential as an informative representation of demographic structures. However, there is a need for novel quantitative approaches that can extend beyond traditional measures like median age or population growth rates to capture the intricate relationships between demographic compositions and their impacts on diverse domains. This gap necessitates the development of new scalar variables that are capable of representing demographic variability in a more robust and actionable manner.
As mentioned earlier, the literature highlights the instrumental role of the population pyramid in policy-making [22, 23, 24]. Policymakers could utilize the proposed variable or mechanism to devise more informed policies. Moreover, the proposed mechanism can be adapted to identify countries or regions within a country, provided sufficient data is available, that share similar socioeconomic standards and enable benchmarking against their disease severity or vulnerability. The ability to correlate demographic structures with health outcomes not only advances our understanding of population dynamics but also lays the groundwork for comparative studies with other demographic and development indicators.
To build on this foundation, we consider related research that investigates similar connections between population metrics and health outcomes. A similar approach was taken by Zhu et al. as they studied the relationship between coronary heart disease and the Human Development Index (HDI)[25]. Additionally, as stated in [13], population density is also considered in investigating the correlation with severity measures. Similarly, numerous socioeconomic indicators exist; however, based on the literature, median age, population density, GDP per capita, HDI, and life expectancy were chosen for comparison in this study due to their strong association with population demographics and their ability to explain variations in health outcomes[26, 27, 28, 29, 30, 31, 32, 33]. Later in the study, the effectiveness of the proposed scalar variables was assessed by comparing their correlation results with these widely used socioeconomic indicators.
Figure 2: Difference in population distribution for similar (a) Median Age, (b) GDP per Capita, (c) Life Expectancy, (d) Population Density, and (e) Human Development Index
Figure 3: Difference in (a) Median Age, (b) GDP per Capita, (c) Life Expectancy, (d) Population Density, and (e) Human Development Index for similar population distributions
In Figure 2, a few examples illustrate the significance of the population pyramid from the other commonly used variables: median age, population density, GDP per capita, HDI, and life expectancy for statistical analysis. In Figure 3, We have provided the population pyramids of the countries with similar indices for the other variables but with disparate demographic distributions and vice versa. The difference suggests that the alternative variables are affected by factors other than the population demographics.
Population pyramids that are similar in shape more often than not correspond to similar median ages. Nonetheless, the converse of that statement does not hold water, as countries with similar median ages might have dramatically different population pyramids, as shown in Figure 2(a). This is because the median age is indifferent to any symmetrical changes in the population pyramid. Accordingly, it is conclusive that countries with closely matching population pyramids have similar demographic profiles, while such a claim is not necessarily valid when using the median age as a parameter.
This study explores the relationship between population demographics and disease severity, aiming to understand how variations in demographic distributions influence national mortality rates for different diseases. By examining these correlations, we seek to uncover patterns that can provide deeper insights into the impact of population structures on mortality causes. Bearing this goal in mind, we introduce two novel scalar variables: PoPDivergence and PoPStat. PoPDivergence is designed to encourage the use of a more informative representation of demographic distributions, while PoPStat quantifies the strength of the correlation between PoPDivergence and a severity measure for a given disease. The methodology used to generate these variables will be properly discussed in the following sections. This study examines the application of these two scalar variables as explanatory variables for various mortality causes, facilitating the correlation of demographic distribution variances with national mortality rates for different diseases. The variable was applied to all-cause mortality data for 2021, specifically focusing on recorded deaths per million. This novel approach aims to statistically quantify the age distribution, or the shape of the population pyramid, in relation to mortality data for various cases of death.
The results highlight how these metrics capture latent information in population pyramids, offering insights into the demographic, social, and cultural factors that influence mortality. The performance of PoPStat was shown to be superior to other demographic and socio-economic indicators, especially for non-communicable diseases (NCDs), and it performed better than traditional development metrics like HDI and GDI in explaining disease mortality patterns. It also provides new perspectives on the Epidemiological Transition Model (ETM)[34] by showing how NCDs align with constrictive population pyramids, while communicable diseases are more associated with expansive pyramids. The findings underscore the utility of PoPDivergence and PoPStat as tools for understanding the complex interplay between population structure and health outcomes, offering potential for future research in other areas of social, behavioral, and economic studies.
Further, the association of the population pyramid to the economy allows allocating funds[35] optimally to critical sectors identified as vulnerable zones based on their demographics now linked more effectively to different causes of death. This signifies the importance of the PoPStat as it can correlate to various socioeconomic and healthcare attributes through the structure of the population pyramid because these socioeconomic and healthcare standards directly impact the death rate. While medical research is about the prevention and eradication of diseases, analysis with an essential tool such as a population pyramid is effective in understanding the factors that affect different rates for mortality causes other than medical.
2 Methods
2.1 Mortality Data
Mortality data was retrieved from the Global Burden of Disease (GBD) Study 2021 conducted by the Institute for Health Metrics and Evaluation (IHME) which includes estimates for 371 diseases and injuries across 204 countries. The causes were systematically organized within a four-tiered hierarchical classification progressively increasing in specificity which are mutually exclusive and collectively exhaustive at each level. Level 1 consisted of three broad aggregate categories (communicable, maternal, neonatal, and nutritional (CMNN) diseases; non-communicable diseases; and injuries). Level 2 disaggregated these categories into 22 cause clusters, while Level 3 further expanded to 175 causes, comprising 132 specific causes and 43 clusters derived from Level 4 causes. For this analysis, causes that had fatal outcomes were examined up to Level 3.
2.2 Population Data
Population data was obtained from the Population Division, Department of Economic and Social Affairs, United Nations as of World Population Prospects 2024. Population pyramids for each country were then constructed for 2021 by organizing the population data by five-year age groups. The data for each age category were then normalized by dividing the population in each group by the total population, effectively representing the proportion of the population for each gender across five-year age intervals.
2.3 Constructing the PoPDivergence
Population pyramids are multidimensional representations of population structure. To establish a scalar parameter that can be correlated with demographic metrics such as mortality rates, we propose PoPDivergence, a distance measure between population pyramids. This approach involves calculating the divergence from a pre-selected reference population pyramid, quantifying the extent to which one population pyramid can be inferred from the reference, and evaluating the deformation between the structural shapes of two population distributions.
PoPDivergence is a Kullback-Leibler (KL) divergence-based metric to measure the relative difference between two probability distributions. Assuming two population distributions $P$ and $Q$, with $P_{(i)}$ and $Q_{(i)}$ representing the proportion of the population in the $i^{th}$ age category, and $Q$ serving as the reference population, we calculate the PoPDivergence as mentioned in equation 1,
$PoPDivergence(ref:Q)=\sum_{i\in I}P_{(i)}log\left(\frac{P_{(i)}}{Q_{(i)}}\right)$ (1)
which quantifies the relative entropy between the two distributions. Specifically, the metric measures the information-theoretic distance from the reference population P to the comparative population Q.
2.4 PoPDivergence Interpretation
Figures 4 and 5 illustrate the quantification of differences between population pyramids using two distinct reference populations; the Central African Republic and Japan. A PoPDivergence of zero indicates no Kullback-Leibler (KL) divergence between population pyramids, which occurs when comparing the reference country with itself. For all other countries, the PoPDivergence reflects the KL divergence between their respective population pyramids and the reference population. When using the Central African Republic as the reference population—a country in the early stages of demographic transition—an increase in PoPDivergence reveals the progression from a progressive to a regressive population pyramid. Conversely, with Japan as the reference population—representing a country in the later stages of demographic transition—an increasing PoPDivergence indicates a shift from a regressive to a progressive population structure. The interpretation of PoPDivergence changes must always be contextualized by the specific reference population.
0
(a) Central African Rep.
0.009
(b) Chad
0.186
(c) India
0.643
(d) Poland
0.664
(e) Malta
Figure 4: PoPDivergence for the reference of Central African Republic
0
(a) Japan
0.026
(b) Finland
0.079
(c) Russia
0.320
(d) Bangladesh
0.518
(e) Kenya
Figure 5: PoPDivergence for the reference of Japan
We consider PoPDivergence as the explanatory variable for disease-specific mortality rates (number of deaths due to a given cause per one million of the population of a given year), to compute the correlation. Since PoPDivergence’s distance computation is not symmetric, we establish the reference population using the objective function below in 2. This maximizes the correlation between cause-specific crude mortality rate in the natural log scale and PoPDivergence.
$\underset{\omega\in\Omega}{\arg{max}}\;\;Cor\left[\ln{S},PoPDivergence(\omega)\right]$ (2)
where $\Omega$ is the set of all countries included in the study, $S$ is the cause-specific crude death rate, and PoPDivergence($\omega$) is the distance between two population pyramids with $\omega$ as the reference country. We used brute force searching to establish the reference population pyramid because there is no closed-form solution to the above optimization problem. It is computationally economical given the discrete nature of the variables involved. This method is what we introduce as reference country tuning, which identifies the population demographic most strongly correlated with a specific disease’s death rates.
2.5 Constructing the PoPStat
PoPStat quantifies the relationship between cause-specific death rates and the structure of population pyramids across countries. To derive this metric, we first determine a reference country for a given cause through reference country tuning. Subsequently, for each country, the PoPDivergence from the reference country is calculated and stored in vector X. The corresponding cause-specific death rates, transformed to the natural logarithm scale, are stored in vector Y. PoPStat is then defined as the Pearson correlation coefficient between X and Y, as calculated using Equation 3, thereby representing the relationship between PoPDivergence (reference) and cause-specific mortality.
$PoPStat(ref)=\frac{\sum_{i=1}^{n}(x_{i}-\bar{x})(y_{i}-\bar{y})}{\sqrt{\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}(y_{i}-\bar{y})^{2}}}$ (3)
where, $x_{i}$ and $y_{i}$ are the PoPDivergence value and natural log of the number of deaths per million for the $i^{th}$ country, respectively, and $x,y$ are the mean values of vectors $X$ and $Y$. In this study, we will consistently present Popstat in the format (PopStat value, reference country, p-value).
2.6 Comparison of PoPStat with other indicators
When comparing the PoPStat metric with other potential indicators, mentioned in the introduction also, we began by identifying the most commonly used demographic indices in the literature, specifically life expectancy, median age, and population density for the countries analyzed. Furthermore, to assess the extent to which population pyramids capture latent information about social and economic determinants of disease mortality, we included Gross Domestic Product (GDP) per capita as a macroeconomic indicator and the Human Development Index (HDI) as a well-established composite development measure for the relevant countries. Next, we calculated the Pearson correlation coefficients between each of these indices and cause-specific death rates to compare their associations of mortality against PoPStat.
3 Results
3.1 Comparison of PoPStat with other indicators
In Figure 6, ischaemic heart disease (IHD) has been used as an example to calculate the correlations of other well-recognized demographic and socio-economic indicators with mortality rate and to compare their performance with PoPStat. As seen in Figure 6(a), the linear variation of the IHD mortality rate has been captured best by the PoPStat with a moderate correlation of - 0.74 (Latvia, p: < 0.001). Meanwhile, the median age also had a moderate correlation of 0.70 (p: <0.001) with mortality due to IHD but it was lower than the PoPStat. As illustrated in Figure 3, PoPStat also showed a better correlation with IHD mortality than the correlations with life expectancy, population density, HDI, and GDP per capita.
(a) PoPStat: ($-0.750$, Latvia, $p:<0.001$)
(b) HDI: $r=0.575$
(c) Median Age: $r=0.708$
(d) Life Expectancy: $r=0.528$
(e) GDP per capita: $r=0.206$
(f) Population Density: $r=-0.039$
Figure 6: PoPStat of Ischemic heart diseases against correlations with other indicators
3.2 PoPStat and disease mortality
3.2.1 Level 1 Disease Categories
Among the three aggregated disease categories in Level 1, the highest PoPStat was observed for non-communicable diseases (NCD) (PoPStat: - 0.84, reference country: Japan, p-value: <0.001) indicating a statistically significant, strong correlation with the underlying shape of the population pyramid across countries. As shown in Figure 7(a), countries with constrictive and more symmetrically shaped population pyramids such as Japan, Italy, Portugal and Greece had higher mortality due to NCDs. The PoPDivergence – which, in this situation, represents the divergence of population from a constrictive pyramid towards an expansive pyramid - explained $71.6\%$ of the variance in global mortality attributed to NCDs.
The lowest PoPStat was observed for injuries (0.29, Singapore, p: <0.001) suggestive of a weak correlation with the demographic structure of the country. Countries like Singapore, Mauritius, Chile and Saint Lucia with constructive pyramids that were largely symmetrical had lower injury-related mortality. However, the deviation from these pyramids towards more progressive shapes explained only 8.4% of the variation in injury-related mortality. As seen in Figure 7(b), the aggregate of communicable, maternal, neonatal and nutritional diseases (CMNN) had a PoPStat of 0.50 (Singapore, p <0.001), indicating a moderate association with the shape of the population pyramid. Countries with more symmetrical and constrictive pyramids like Singapore, Mauritius, Chile and Saint Lucia saw lower CMNN death rates. The divergence from these regressive population structures toward progressive pyramids accounted for nearly one-fourth of the change in mortality attributed to CMNN causes.
(a) Non-Communicable Diseases
(b) Communicable, maternal, neonatal, and nutritional diseases
(c) Cardiovascular Diseases
(d) Neglected Tropical Diseases
Figure 7: Main caption describing the entire figure
3.2.2 Level 2 Disease Categories
Non-Communicable Diseases
Among NCDs, the largest PoPStat (0.90, p: <0.001) was noted for neurological diseases, with Zambia as the reference population pyramid. Similar countries like Burkina Faso, Cameroon and Yemen - with expansive, predominantly symmetrical population pyramids - had a lower burden of deaths due to neurological diseases. Shifting away from these early stages of the demographic transition explained 81.4% of the variation in mortality assigned to neurological causes. The next highest PoPStat was reported for neoplasms (-0.90, Japan, p: <0.001). Populations with symmetrical, constrictive pyramids - for example, Japan, Italy, Portugal and Greece - were associated with higher cancer mortality rates in the world. 80.4% of the change in cancer mortality could be explained by the PoPDivergence from a constrictive reference pyramid. Cardiovascular diseases (CVD) had a strong PoPStat of – 0.73 (p: <0.001), with Japan as the reference population. Figure 7(c) illustrates how countries with symmetrical and constrictive population pyramids identified before had more death rates due to cardiovascular diseases as well. The variance in cardiovascular mortality explained by the divergence of populations from constrictive pyramids towards expansive pyramids was 53.4%. Meanwhile, mental disorders also had a strong PoPStat of – 0.70 (Japan, p: <0.001). Similar countries with symmetrically constrictive population pyramids shared a higher burden and shifting from them toward more progressive shapes explained nearly half of the change in deaths due to mental illnesses
Musculoskeletal disorders had a moderate PoPStat of – 0.64 (Japan, p: <0.001). Deaths due to musculoskeletal disorders were also associated with similar constrictive population pyramids and divergence from them explained 41.8% of the change in mortality. The PoPStat for chronic respiratory diseases was -0.53 (p<0.001) with France as the reference country. Constrictive population pyramids like Finland, Belgium and the UK which were less inverted than former examples saw higher death rates due to chronic respiratory diseases. Deviation from this demographic structure explained 28.1 of the change in fatalities due to chronic respiratory illnesses. Substance use disorders also showed a moderate PoPStat of -0.53 (p: < 0.001) with Japan as the reference. Countries like Italy, Portugal and Greece, characterized by symmetrical constrictive pyramids, experienced higher mortality. The divergence from these pyramids explained 28.3% of the variance in substance-related deaths. Digestive diseases had a moderate PoPStat of 0.50 (p: <0.001) with Qatar as the reference population pyramid. Countries with constrictive pyramids that were asymmetrical or male-predominated – like the United Arab Emirates, Oman and Maldives – had lower mortality due to digestive diseases. PoPDivergence in relation to such pyramids explained nearly one-fourth of the variation in deaths caused by digestive disorders. The mortality due to other NCDs category (0.49, Singapore, p: <0.001) was associated with symmetrical and constrictive population pyramids. PoPDivergence from these pyramids accounted for 23.4% of the change in deaths due to other NCDs. Meanwhile, diabetes and kidney diseases maintained the lowest PoPStat of -0.41 (p: <0.001), with Argentina becoming the reference population. Similar countries with late-expansive population pyramids – Costa Rica, Sri Lanka and St. Vincent and Grenadines – had higher mortality attributed to diabetic and kidney diseases. Nonetheless, the change in diabetic deaths explained by the departures from a late-expansive pyramid was 16.6%.
Communicable Diseases
As visualized in Figure 7(d), the highest PoPStat among communicable diseases was observed for neglected tropical diseases (NTD) (0.83, Malta, p: < 0.001), indicating a very strong association between these diseases and the demographic structure. Countries with symmetrical, constrictive population pyramids such as Malta, Poland, Cyprus and Slovakia experienced lower mortality from NTDs. The PoPDivergence—reflecting movement from constrictive pyramids toward expansive pyramids—explained 69.7% of the variance in mortality related to these conditions. A similar pattern emerged for other infectious diseases which had a very strong PoPStat of 0.82 (p: < 0.001). Malta again served as the reference population with countries like Poland, Cyprus and Slovakia exhibiting lower mortality rates due to other infections. The shift from these symmetrically constrictive pyramids towards expansive shapes amounted to 68.4% of the variation in mortality from other infectious diseases.
Enteric infections demonstrated a strong PoPStat of 0.73 (p: < 0.001). Similar countries like Malta, characterized by symmetrical and constrictive pyramids, had lower mortality rates. The divergence from these structures towards early-stage pyramids explained 54.0% of the variance in enteric infection-related mortality. HIV and sexually transmitted infections (STIs) showed a moderate PoPStat of -0.64 (p: < 0.001) with the Central African Republic as the reference country. Countries with symmetrical and expansive demographic pyramids such as Niger, Somalia and Chad exhibited higher mortality rates. Deviations away from these early-stage pyramids explained 41.5% of the change in mortality. Respiratory infections and tuberculosis had the lowest PoPStat among communicable diseases (0.14, Qatar, p: < 0.001), suggesting a weak association with demographic structures. Asymmetrically constrictive pyramids like Qatar, UAE, Bahrain and Maldives had a lower burden of deaths. However, divergence from these pyramids only accounted for 2.0% of the variance in mortality.
Maternal, neonatal and nutritional diseases
Maternal and neonatal conditions exhibited a very strong PoPStat 0f -0.91 (p < 0.001) with the Central African Republic as the reference. Countries such as Niger, Somalia and Chad which shared expansive population pyramids saw higher death rates. Deviations from these structures explained 83.6% of the variance in maternal and neonatal mortality. Meanwhile, nutritional deficiencies showed a moderate PoPStat of 0.56 (Cyprus, p: < 0.001). Countries like Luxembourg, Chile and Iceland - having symmetrical, constrictive pyramids- exhibited lower mortality rates. Transitioning away from these pyramids towards expansive ones explained 32.3% of the change in mortality.
Injuries
Transport injuries (TI) (-0.62, p: < 0.001) indicated a moderate association with the underlying demography across countries. Burundi served as the reference population with countries like Mali, Burkina Faso and Mozambique showing similar symmetrically expansive pyramids and higher mortality. Movement away from these population pyramids accounted for 38.5% of the variance in mortality. Unintentional injuries (UI) displayed a moderate PoPStat of 0.52 (p: < 0.001). Kuwait was the reference population and countries with both expansive and stationary pyramids like Brunei, Iran and Bahrain showed higher mortality rates. The PoPDivergence explained 27.4% of the variation. Self-harm and interpersonal violence had a very weak PoPStat of 0.18 (p: < 0.001). Qatar was the reference and similar asymmetrically constrictive pyramids like Bahrain and the UAE had lower mortality rates. The shift in demographic structures accounted for only 3.2% of the variation of mortality due to self-harm and interpersonal violence.
Table 1: PoPStat and reference populations for selected Level 3 diseases
Disease Reference PoPStat $r^{2}$ p-value Closest population 2nd Closest population 3rd Closest population
Ischemic heart disease Latvia -0.751 0.562 <0.001 Estonia Lithuania Slovenia
Stroke France -0.497 0.247 <0.001 Belgium United Kingdom Finland
Pulmonary Arterial Hypertension Oman 0.375 0.14 <0.001 Saudi Arabia Maldives Bahrain
Chronic obstructive pulmonary disease France -0.621 0.386 <0.001 Belgium United Kingdom Finland
Asthma Central African Republic -0.594 0.352 <0.001 Niger Somalia Chad
Breast cancer Japan -0.815 0.664 <0.001 Italy Portugal Greece
Colon and rectum cancer Japan -0.922 0.85 <0.001 Italy Portugal Greece
Cervical cancer Tonga -0.257 0.066 <0.001 Samoa Syria Comoros
Prostate cancer Japan -0.716 0.513 <0.001 Italy Portugal Greece
Liver cancer Oman 0.499 0.249 <0.001 Saudi Arabia Maldives Bahrain
Cirrhosis and other chronic liver diseases Qatar 0.162 0.026 0.03 United Arab Emirates Bahrain Maldives
Inflammatory bowel disease Oman 0.634 0.401 <0.001 Saudi Arabia Maldives Bahrain
Alzheimer’s disease and other dementias Japan -0.925 0.855 <0.001 Italy Portugal Greece
Parkinson’s disease Japan -0.936 0.877 <0.001 Italy Portugal Greece
Alcohol use disorders Oman 0.438 0.192 <0.001 Saudi Arabia Maldives Bahrain
Diabetes mellitus Sri Lanka -0.355 0.126 <0.001 Saint Vincent and the Grenadines Brazil Argentina
Chronic kidney disease Uruguay -0.411 0.169 <0.001 New Zealand Iceland Australia
Rheumatoid arthritis Puerto Rico -0.713 0.509 <0.001 Portugal Croatia Finland
Maternal disorders Central African Republic -0.909 0.827 <0.001 Niger Somalia Chad
Neonatal disorders Central African Republic -0.91 0.827 <0.001 Niger Somalia Chad
Self-harm United Arab Emirates 0.443 0.197 <0.001 Qatar Bahrain Maldives
Interpersonal violence Tonga -0.472 0.223 <0.001 Samoa Syria Comoros
HIV/AIDS Central African Republic -0.549 0.302 <0.001 Niger Somalia Chad
Tuberculosis Malta 0.759 0.576 <0.001 Poland Cyprus Slovakia
Dengue Jamaica -0.578 0.334 <0.001 Colombia Saint Lucia Brunei
Protein-energy malnutrition Malta 0.561 0.315 <0.001 Poland Cyprus Slovakia
3.2.3 Level 3 Disease Categories
Table 1, presents PoPStat related to selected Level 3 disease categories, along with their corresponding reference countries, countries adjacent to reference populations and other important statistical measures.
4 Discussion
4.1 PoPDivergence and PoPStat
In this study, we devised two novel metrics, PoPDivergence and PoPStat. PoPDivergence measures the difference between the shape of a country’s population pyramid and that of a reference population pyramid. PoPStat is the correlation between PoPDivergence and the (natural log form of) death rate for a disease of interest. We also designed an optimization mechanism to select a reference population that invariably maximizes this correlation. PoPStat, in other words, is a measure that tells you how tightly a specific disease is associated with the population pyramid. To the best of our knowledge, this is the first study to have introduced a metric that quantifies this association. PoPStat also provides information on the nature of population structures most or least compatible with high mortality from that specific disease.
Meaningfully summarizing the multi-dimensional population structure into a scalar variable allows us to explore its relationship with various outcomes. Though the current study has focused on disease mortality, the correlation with PoPDivergence can be assessed for any social, behavioural or economic phenomena of choice. On the exposure end, a range of causal information is captured by the shape change in the population pyramid. Firstly, the shift in sex and age composition between pyramids directly affects gender- and age-specific mortality rates of a specific cause. Secondly, despite being a direct function of the rate of birth, deaths and migration, the population pyramid also embodies multiple biological, socio-economic, behavioral and cultural attributes that directly or indirectly compound disease-specific mortality. Thirdly, the shape of population pyramids also carries a wider representation of the geographic or regional characteristics that may impact global mortality patterns.
One major challenge to studying the dynamics of population change has been the resource-intensiveness of analyzing data not only across geographies but also across generations. But, each country and region is at a unique point in its demographic transition as reflected by their population pyramids. Therefore, PoPDivergence allows us to capture the longitudinal axis of global population change within a single cross-section in time. The observation that the reference population mostly happens to be at either end of the demographic transition pyramid is also noteworthy. For any given disease category, with increasing PoPDivergence, we see a transition of population pyramids similar to that described by the demographic transition model without any explicit instructions to follow this path. So, any monotonic metric linked to PoPDivergence will allow us to empirically visualize the epidemiological, social, and economic transitions that happen parallel to the demographic transition.
4.2 Performance of PoPStat with other indicators
Examining IHD as an illustrative example, the association of IHD mortality with PoPDivergence - represented by the PoPStat - was superior to the correlations of other demographic indicators included in this study. This pattern remained true for all the relevant causes and cause clusters across the three levels (Supplemental File 1). This indicates that other scalar variables such as median age and life expectancy may condense lesser information about the demography and its health compared to the PoPDivergence. Therefore, PopDivergance becomes a high-quality dimension reduction of the information-rich population pyramid. For example, in relation to Figure 2(a), PoPDivergence captures the shape differences of the two population pyramids that share a singular median age. However, it must be noted that the correlation between median age and disease mortality remained equally strong for a majority of causes (Supplemental File 1). So, future research can explore if the ability of these two indicators to associate with the underlying demographic structures of the demographic transition remains true for areas other than epidemiology.
PoPStat also performed better than the disease mortality correlations of other development-related indicators. It was noteworthy that PoPStat had a better association with the social determinants of IHD mortality than HDI which was a latent variable composite of life expectancy, years of schooling and Gross National Income (GNI) per capita. It was also superior to the correlation between IHD mortality and GDI per capita taken alone. However, it has to be noted that the relationship between health and the national income of a country is likely curvilinear - similar to Preston’s curve[36]. Pearson’s correlation may not do justice in analyzing such an association. Nevertheless, these findings demonstrate the versatility of PoPDivergence as a measure of the population structure during disease burden assessments. It does not only associate more strongly with demographic determinants of mortality compared to other population descriptors but also aligns more closely with socio-economic determinants of mortality than conventional development metrics.
4.3 Level 1 Disease categories
The higher PoPStat of NCDs indicated that they are more strongly associated with the underlying population pyramid of a country and its latent constructs than CMNN diseases. This disparity may have arisen due to intrinsic disease attributes, such as chronic diseases being selective for older ages while infections cross-cut age strata. It may have also been due to the pooling of heterogeneous entities like communicable, maternal, neonatal and nutritional causes, which may have diluted the measure of association. The lowest PoPStat was noted for injuries. The largest contribution to injuries is made from unintentional injuries such as falls, drowning, fire and natural disasters whose risk may similarly spread across several age cohorts[37]. This may have resulted in the poor association of PoPDivergence with external causes of death compared to natural causes of death that originate from disease processes of ageing. Under-reporting of the true burden of injuries by national surveillance systems may have also contributed to the poor correlation with the countries’ demographics.
The Epidemiological Transition Model introduced by Abdel Omar in 1971 expanded Landry’s DTM by incorporating disease and mortality patterns[34]. One major proposition in ETM was that parallel to the demographic transition, infectious diseases would be gradually replaced by degenerative and man-made diseases as the leading cause of death. However, the validity of ETM has been heavily contested in literature due to the linearity in its transition and the many deviations observed. Santhosa et al. (2014) in their review of literature over four decades identified a discordance between the theory of ETM and empirical evidence[38]. Therefore, the observation of NCD mortality being associated with constrictive population pyramids whereas the burden of deaths due to CMNN diseases being linked to expansive pyramids critically adds to the current discourse on ETM. Our findings indicate that when the demographic stages of all the countries are captured within a single timeframe, the propositions of ETM may hold water even after half a century. However, it must be noted that PoPDivergence cannot assess the pace of these transitions and cannot identify which ET model: classical, accelerated or delayed, fits the contemporary global picture best.
Considering raised NCD mortality among regressive population structures, in 2019, 18777 Years of Life were Lost (YLL) due to NCDs in very high Human Development Index (HDI) countries compared to 7444 YLL in very low HDI countries[39]. Omar identified socio-economic, political, cultural, ecological and health system determinants for this epidemiological transition[34]. First, the pathological processes underlying the most fatal NCDs - cardiovascular diseases and cancers - are closely associated with the age structure, making them highly prevalent among aging populations[40]. Next, the crude prevalence and attributable burden of primary risk factors for NCDs such as physical inactivity and alcohol consumption; and intermediate risk factors like hypertension were also greater in the GBD regions: High Income North America, High Income Asia Pacific, Western, Central and Eastern Europe, with predominantly constrictive population pyramids[41, 42, 43]. Even after adjusting for age, the prevalence and attributable burden of risk factors like tobacco smoking, alcohol use, dietary risks and obesity remained elevated for most of these GBD regions[44, 45, 46]. This indicates that the strong PoPStat for NCDs may be due to capturing latent information within population pyramids that extend beyond age and sex dynamics.
Considering environmental exposures associated with the urbanization and industrialization complex, these regions with constrictive pyramids interestingly had a lower attributable burden of disease due to air pollution. However, the attributable cancer mortality due to occupational carcinogens was contrastingly higher than in other populations[47, 48]. These are the likely risk exposures that have led to a higher NCD burden in regressive population pyramids. Finally, health system factors also explain the burden of NCDs in constrictive population pyramids. Advanced health technologies and higher per capita health expenditures may have contributed to longevit y and the consequent chronic health problems of those aging populations.
The relationship between communicable diseases and expansive population pyramids also operates on multiple levels. First, communicable diseases with the highest mortality - for example, enteric infections - mainly affect younger populations. Therefore, the age structure of broad-based population pyramids has a direct impact on infectious disease mortality[49]. Next, the burden of malaria, tuberculosis, HIV/AIDs and neglected tropical diseases is primarily held by countries with expansive population pyramids in Africa, Asia and Latin America[50]. This may be due to multiple socio-economic determinants such as poverty and colonial past that are shared by them or due to common ecological characteristics such as temperate climate that facilitate the survival of specific pathogens and vectors[51].
The association between injuries and expansive population pyramids has been previously documented in the literature. For example, in 2021, age-standardized Disability Adjusted Life Years (DALY) rates for injuries among males were greater in the low Social Development Index (SDI) countries (5958.8 per 100,000) compared to the high SDI countries (3050.8 per 100,000)[50]. This may be attributed to low-quality infrastructure, increased occupational hazards, lack of safety standards, limited injury prevention and disaster preparedness efforts in countries at the early stages of demographic transition[37].
4.4 Level 2 Disease Categories
4.4.1 Non Communicable Diseases
Considering Level 2 NCD categories, the highest PoPStat was seen among neurological disorders. This is understandable as the largest mortality due to neurological disorders is caused by strokes and degenerative conditions like Alzheimer’s disease which are mostly associated with older age and thereby the demographic structure[52]. The PoPStat of neurological disorders was even greater than cancers which tend to manifest in children and young ages despite their high overall burden. However, following age standardization, regions with expansive population pyramids also showed an increased number of DALYs indicating that the association of neurological disorders to population pyramids was largely mediated by age[52].
Association of mental disorders and constrictive population pyramids aligned with the existing knowledge base as the highest number of DALYs due to mental diseases were observed among populations in North America, Australia and Western Europe that were in the latter stages of the demographic transition[53]. Substance use disorders being linked to constrictive population pyramids was also predictable given the most used substance was alcohol which had the largest crude prevalence in Western and East European regions and North America after the Asian region[45]. Since alcohol and drug use was not limited to an age stratum[45], in relation to substance use disorders, PoPStat did not only correlate with the age structure but also reflected socio-cultural determinants that these populations may have acquired during their demographic transition.
Chronic obstructive pulmonary disease (COPD) is the main cause of chronic respiratory disease mortality worldwide[54]. Chronic respiratory diseases being connected to less-inverted population pyramids is explainable on this basis where a higher burden of COPD deaths was observed among Western European population pyramids with comparable shapes[54]. However, when adjusted for age, the highest death rates of chronic respiratory diseases were seen among African and Asian regions with expansive population pyramids. This indicates that - similar to neurological disorders - the association of chronic respiratory diseases with the population pyramid was mediated by the age structure. The reduced burden of digestive disease mortality among asymmetrically constrive population pyramids in the Middle East was noteworthy. Given the second largest cause of digestive disease deaths is cirrhosis, this may be attributed to the cultural characteristics of such population pyramids where alcohol consumption is limited[55, 45].
Diabetes was the sole exception to the relationship between NCD mortality and regressive population pyramids observed in this study. Higher diabetic mortality among expansive population pyramids was in line with the current evidence where the death rates were highest in the GBD super regions of Latin America and the Caribbean, North Africa and the Middle East, Sub-Saharan Africa and South Asia[56]. This deviation could be due to a myriad of causes: high BMI or obesity in regions such as North Africa and the Middle East, genetic predispositions among regions such as South Asia and determinants such as poverty, limited access to healthy food and lack of comprehensive diabetic care that are spread across these regions[56]. It is also important to note that diabetes carried the lowest PoPStat among NCDs. This suggests that the contribution of some of the above factors may not be captured by the population pyramid. However, compared to other demographic and developmental indicators, PoPStat still had the best correlation with diabetic mortality (Supplemental File 1).
4.4.2 Communicable Diseases
NTDs had the highest PoPStat among the Level 2 communicable diseases and the majority of the burden was shared by expansive population pyramids. NTDs are a heterogeneous group of parasitic, bacterial and viral infectious diseases that commonly occur in tropical regions, predominantly among low and middle-income countries of Asia, Africa and Latin America. However, between 1990 and 2019 the most pronounced increase in NTD incidence was observed among the middle and high SDI regions[57]. Our findings indicate that the spread of NTDs to developed populations away from the equator has not diluted their association with progressive population pyramids in the tropics. This may be due to haphazard urbanization and poor waste management of growing populations that support vector breeding, limited public health infrastructure including sanitation and other geographical and environmental attributes shared by these populations[57]. However, the correlations of NTD mortality with UHC and HDI were seen to be higher than PoPStat (Supplemental File 1). This suggests that the impact of factors such as primary health care delivery and education may be larger on NTDs than direct and indirect demographic influence.
Other infectious diseases had a strong PoPStat, with the major mortality seen among expansive population pyramids. This was partly due to the largest burden of other infections being caused by meningitis - a condition frequently encountered among expansive population pyramids in the African region[52]. However, the correlation of the mortality of other infectious diseases with UHC was also higher than PoPStat implying that health system factors may play a greater role in communicable disease mortality than demographic variables. HIV and other STIs had a moderate PoPStat and the burden of deaths reduced in the shift from progressive to regressive demographic structures. HIV could be the main culprit behind this finding as it made the largest contribution to this Level 2 category with the highest mortality rates seen among expansive populations in the Sub-Saharan African region[58]. Respiratory infections had the lowest PoPStat among communicable diseases. This was likely due to respiratory infections and their causative pathogens affecting both sexes in all age strata and thereby not being discriminated by the sex-age structure. However, further investigations are needed to see why respiratory infections were low among asymmetrically constrictive pyramids such as Middle Eastern populations. This may have been due to the dry climate, lack of seasonal changes and other regional characteristics shared by them than due to an intrinsic feature of the unique shape.
4.4.3 Maternal, neonatal and nutritional diseases
Among other CMNN categories, maternal and neonatal conditions showed a very strong PoPStat. This tight association with the demography may seem self-evident as a higher number of mothers and neonates in broad-based populations may result in a higher burden of maternal and neonatal deaths. Nevertheless, even after adjusting for age, the greatest DALY rates due to most neonatal causes were from low SDI countries in Asia and Africa and the highest DALY rates due to all maternal causes were from Sub-Saharan Africa. Meanwhile, the lowest age-standardized DALY rates due to all neonatal and maternal causes were from high-income Asia Pacific, North America and Australasia regions[59]. So, the reduction in mortality due to maternal and neonatal causes when transitioning from progressive to regressive population structures is not merely due to changes in the pyramid base but is also influenced by socio-economic and health determinants acquired during the process. Nutritional deficiencies carrying a moderate PoPStat was interesting given death rates due to malnutrition being strongly polarized between low SDI (11.5 per 100000) and high SDI (0.6 per 100000) regions in a manner similar to maternal and neonatal causes[46]. WHO’s Eastern Mediterranean region showing a mixed picture with higher death rates due to both malnutrition and obesity could have dampened the PoPStat of nutritional-related mortality[46].
4.4.4 Injuries
Higher mortality of transport injuries among expansive population pyramids tallied with the literature. Globally, the highest DALYs due to TIs were reported between 20 to 40 years old indicating a direct age-related phenomenon[50]. After age standardization, the bulk of DALY rates due to TIs were still noted among low SDI regions with rising trends in Western Sub-Saharan African, Southern Latin American, South and East Asian populations [60]. This meant that in addition to the accident-prone age groups, environmental determinants such as unsafe roads, cultural determinants like the use of motorbikes and tuk-tuks and political determinants such as limited enforcement of traffic laws may be layered within the associated population pyramids. Interestingly, it has been recognized that TI incidence had a positive correlation with SDI in most of the above regions[60]. This suggests that certain dimensions of the development transition such as urbanization and motorization may themselves increase TI mortality. This could have also attenuated the original PoPStat for transport injuries.
An interesting PoPDivergence emerged when examining the mortality of unintentional injuries. The reference population of Kuwait was relatively in the middle of the demographic transition process while having a lower burden of UI mortality. As KL divergence happens away from this reference it would lead to either predominantly young or predominantly old populations. Therefore, based on the optimization followed in this study, higher UI mortality appears to be associated with populations at either extremities in age. This made sense given the largest burden of UI was due to falls and drowning where both children and older persons carried the highest risk[50].
Self-harm and interpersonal violence demonstrated a very weak PoPStat. This was noteworthy given the increase in self-harm during the development transition has been theorized early in sociology. Emilie Durkheim who studied suicides in the latter part of the 19th century proposed that self-harm was governed by the changes of social integration during the shift of societies from mechanistic to organic solidarity[61]. However, our findings indicate that self-harm and violence may not make such a strong discrimination between developing and developed populations. Though both conditions shared a higher mortality burden among expansive population pyramids than constrictive ones, 98% of the variance in mortality could not be explained by the said transition. This is also empirically supported by GBD studies where high-SDI countries had the second highest age-adjusted mortality rate (11.1 per 100000) after low-middle SDI countries (11.9 per 100000) attributed to self-harm[62]. Though risk factors such as intimate partner violence, high temperatures and socio-political unrest may explain interpersonal violence among expansive population pyramids in the tropical regions, self-harm seems to affect all populations irrespective of their position in the demographic change[62].